# Stability analysis through folds: An end-loaded elastica with a lever arm

### v3

@@ 1 Introduction
The study of physical problems using the variational framework has a rich history. In this setting, the equilibria are characterized as critical points of an energy-like functional. Typically, these problems are expressed in terms of a control parameter, generating a parameter family of equilibria. Indeed, the equilibria are computed by a slow scan of a parameter from a known solution using a continuation technique [1]. However, not all equilibria are stable and physically relevant; they must correspond to the local minima of the functional. Hence, it is crucial to determine whether the critical points are local minima (or stable) and track their changes. Generally, the stability of parameter-dependent equilibria changes at singularities, such as a fold or bifurcation in the parameter space [2, 3, 4]. However, it remains unclear whether a fold in the unstable equilibrium results in a stable equilibrium or a higher-order unstable equilibrium. This direction of stability transition can be encoded through a particular projection of solutions, using distinguished bifurcation diagrams, and the stability can be assessed with a little observation and without any additional analysis. This idea of topologically exploiting the variational structure for stability examination was first employed in Astrophysics [5, 6]. Distinguished bifurcation diagrams were first developed for finite-dimensional mechanical systems [7] that are subsequently extended to infinite-dimensional systems [8, 9]. Since then, several versions of these diagrams have appeared in investigations of elastic problems [10, 11]. In most studies, the diagrams are limited to cases with fixed-fixed ends or cases where the varying parameter appears within the integrand or as a Lagrange multiplier. In this study, we generalize the distinguished bifurcation diagrams to variational problems subject to fixed-free ends, when the bifurcation parameter appears in the boundary conditions. Consequently, two relevant cases arise: when the parameter appears at the fixed end and when it appears at the free end. For this analysis, we introduce the notion of a stability index that is equal to the dimension of the subspace over which the associated second variation operator is negative. The critical points with index zero are local minima and stable. Alternatively, the energy of the system can be used to determine the direction of stability [12, 13]. The lower branch of the fold corresponds to the lower-energy state.
@@
The developed methodology is applied to test the stability of elastica under an end load through a lever arm. Cantilever structures with a lever arm are common in nature and engineering—for example, in plants supporting fruits or in load-bearing flexible systems. In recent years, nonlinear rod models [14] have been actively employed in the fields of biophysics [15], computer graphics [16, 17] and soft robot manipulators [18, 19]. Elastica, a simple planar version of this rod model, has equally received attention. A flexible elastic structure subjected to loads or constraints may exhibit multiple equilibria. Hence, it is crucial to determine their stability, as only stable structures are physically viable. Unstable equilibria may exist under specific constraints or boundary conditions and transitions to the available stable equilibrium under disturbances. Indeed, the presence of instability also hints at snap-back phenomena, a characteristic of multi-stable systems [20]. Numerous studies have been conducted on the stability of elastic rods, which involve either the Jacobi test [21, 22, 23, 24, 25, 26] or Hessian eigenvalue determination [27, 28]. Alternative techniques based on bifurcation analysis are also available [29]. In this article, we assess the stability without performing any such rigorous computations. But, instead, conclusions are drawn by qualitatively examining the plot of appropriate projection of the equilibria against the parameter.
@@
Advances in material science produced highly deformable alloys and polymers, accelerating the development of soft robotics. Generally, soft robots employ very flexible structures to generate compliant mechanisms. A flexible elastic structure under load can exhibit multiple equilibria, where one equilibrium transitions to another through snap-back instability. These bistable or multi-stable systems exhibit large rapid displacements for small stimuli when operated around the regions of instabilities. As the system surpasses the initial energy barrier, it transitions to a new state, accompanied by a release of energy. Some of nature’s quickest mechanisms seen in Mantis shrimp [30], Venus fly trap [31], and Hummingbird beak [32] can be attributed to this snap-back instability. These mechanisms are widely used in the design of high-performance soft actuators [33], soft robots [34], deployable and morphing structures [35] and haptic devices [36]. Inspired by these expanding applications, we aim to investigate elastica with a lever arm and the dependence of its stability on several parameters. These findings could offer invaluable insights for designing innovative mechanisms. Particularly, this study signifies the distinguished bifurcation diagrams for the analysis of multi-stable systems in engineering applications.
@@
This article is organized as follows. The section 2 introduces the unconstrained calculus of variation problem with a parameter in the boundary conditions and describes the associated distinguished bifurcation diagrams. Section 3 formulates the elastica with an end load acting through a lever arm. A bifurcation analysis of resulting equilibria is also performed. In section 4, we present several examples investigating the impact of varying parameters on stability using distinguished bifurcation diagrams. Finally, section 5 provides a summary and discussion of the results.
@@ 2 The Unconstrained Variational Problem
In this section, we extend the theory of distinguished bifurcation diagrams developed by [8] and [9] to variational problems subject to fixed-free ends. We consider both cases, where the bifurcation parameter appears at the fixed end and when it appears at the free end. In section 2.1, we introduce the problem and the relevant notation. The distinguished diagram at a simple fold when the parameter appears at the fixed end is developed in section 2.2 and when the parameter appears at the free end is developed in section 2.3.
@@ 2.1 Problem Formulation
Let $\boldsymbol{\zeta}:s\rightarrow\mathbb{R}^{p}$ be a vector-valued function with a continuous derivative with respect to the independent parameter $s\in[0,l]$. Consider an unconstrained calculus of variations problem of the form
[…]
@@
while the other end $s=l$ is held free. The scalar function $\mathcal{L}$ has continuous second derivatives with respect to all its arguments and is convex in its second argument. On the other hand, the scalar function $B$ has continuous second derivatives with respect to its first argument and has continuous first derivatives with respect to the other arguments. The system can be influenced by the parameters $\xi_{0}$ and $\xi_{l}$ at the boundaries $s=0$ and $s=l$ respectively. In the current study, we vary either $\xi_{0}$ or $\xi_{l}$, but not both at a time. Maddocks [8] generated the bifurcation diagrams for problems when the parameter appears in the integrand and with homogeneous boundary conditions, which were later extended to the case when the parameter appears in the fixed end in a problem with non-homogeneous boundary conditions [9]. However, these studies are confined to problems with fixed-fixed ends. Here, we extend the analysis to the case where one of the fixed ends is set free. The first-order condition yields the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, $i=0$ or $l$, of (1) as solutions to the well-known Euler-Lagrange equations
@@
These critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, $i=0$ or $l$ exist as a continuous curve of solutions along a branch for some pseudo-arclength parameterization $\tau$ [1] satisfying the property
$\big{|}\dot{\boldsymbol{\zeta}}(s,\tau)\big{|}^{2}+\big{|}\dot{\xi_{i}}(\tau)\big{|}^{2}=1,\qquad i=0\text{ or }l.$
The notation $\dot{\boldsymbol{\zeta}}(s,\tau)$ denotes the derivative with respect to pseudo-arclength $\tau$, whereas ${\boldsymbol{\zeta}}^{\prime}(s,\tau)$ denotes the derivative with respect to independent parameter $s$. The second variation $\delta^{2}J$ evaluated at the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, ($i=0$ or $l$) must satisfy the condition [37]:
[…]
@@
For brevity, the dependence of $\mathbf{P},\mathbf{C}$ and $\mathbf{Q}$ on $s$ is not explicitly shown. The matrices $\mathbf{P}$,$\mathbf{Q}$ and $\mathbf{B}$ are symmetric, whereas the matrix $\mathbf{C}$ may not be. Here, $\mathbf{h}$ is a variation in the solution that satisfy the boundary condition:
@@
where $\mathcal{S}$ is the second-order self-adjoint differential operator
@@
[…]
@@
The operator $\mathcal{S}$ is assumed to have eigenvalues and eigenfunctions that smoothly depend on the pseudo-arclength $\tau$. Therefore, it is not sufficient to merely determine the sign of the eigenvalue; it is also necessary to account for the additional boundary contribution. The index for the case of fixed-fixed ends is shown to be equal to the number of conjugate points (for more details refer [21]). However, we will not pursue our analysis in that direction. Instead, we track changes in the index at critical points.
@@
$\langle\mathcal{S}\mathbf{u},\mathbf{v}\rangle=\langle\mathbf{u},\mathcal{S}\mathbf{v}\rangle+\Big{[}\left(\mathbf{P}\mathbf{v}^{\prime}+\mathbf{C}^{T}\mathbf{v}\right)\cdot\mathbf{v}\Big{]}^{l}_{0}-\Big{[}\left(\mathbf{P}\mathbf{u}^{\prime}+\mathbf{C}^{T}\mathbf{u}\right)\cdot\mathbf{v}\Big{]}^{l}_{0},$ (13)
through successive applications of integration by parts. Therefore, if boundary terms vanish $\mathcal{S}$ is a self-adjoint operator.
Parameter-dependent variational problems generate a family of extremals, and in the case of a single parameter $\xi$, a curve of solutions is obtained. Sometimes, these solutions are characterized by folds- points at which the curve slopes vertically. In bifurcation terminology, these points are also referred to as saddle-node bifurcation or turning points. On differentiating the Euler-Lagrange equations given by (3) and the non-linear boundary conditions (4) with respect to pseudo-arclength $\tau$, we obtain the following expressions for the two cases of varying parameters
@@
Both of these boundary value problems correspond to the eigenvalue problem (11) with zero eigenvalue whenever $\dot{\xi}_{0}=0$ or $\dot{\xi}_{l}=0$, with the eigenvector $\mathbf{h}$ coinciding with $\dot{\boldsymbol{\zeta}}$. The boundary condition at $s=l$ also removes the boundary term in (12), fetching a zero second-variation quadratic functional. We focus on the behavior of the motion of this functional around these special points - folds. A simple fold further satisfies:
$\ddot{\xi}_{i}\neq 0,\qquad i=0\text{ or }l$ (15a)
$\frac{d}{d\tau}J_{\xi_{i}}\neq 0\qquad i=0\text{ or }l,$ (15b)
and $\text{and }\mu=0\text{ is a simple eigenvalue of }\mathcal{S}.$ (15c)
An eigenvalue is called a simple eigenvalue if its algebraic multiplicity is one. The meaning of (15b) becomes clearer towards the end. We analyze the behavior of the second-variational functional in the vicinity of these folds, which involves simultaneously monitoring the eignevalues and the boundary contribution from the corresponding eigenfunctions. We assume that changes in stability occur exclusively in the vicinity of critical points, such as folds or bifurcation points [2].
@@
$\frac{d}{d\tau}\delta^{2}J[\mathbf{h}]=\langle\dot{\left(\mathcal{S}\mathbf{h}\right)},\mathbf{h}\rangle+\dot{\left(\mathbf{P}\mathbf{h}^{\prime}+\mathbf{C}^{T}\mathbf{h}+\mathbf{B}\mathbf{h}\right)}\cdot\mathbf{h}\bigg{|}_{s=l}+\langle\mathcal{S}\mathbf{h},\dot{\mathbf{h}}\rangle+\left(\mathbf{P}\mathbf{h}^{\prime}+\mathbf{C}^{T}\mathbf{h}+\mathbf{B}\mathbf{h}\right)\cdot\dot{\mathbf{h}}\bigg{|}_{s=l},$ (16)
@@
Since $\dot{\boldsymbol{\zeta}}$ is the zero eigenvector of $\mathcal{S}$, we have $\mathcal{S}\dot{\boldsymbol{\zeta}}=\mathbf{0}$ implying $\langle\dot{\mathbf{h}},\mathcal{S}\dot{\boldsymbol{\zeta}}\rangle=0$. On applying the boundary conditions denoted by (14) and (17) at a fold $\dot{\xi_{i}}$ ($i=0$ or $l$), the above expression simplifies to
@@
$\ddot{\boldsymbol{\zeta}}(l)$ $\equiv\mathbf{P}\ddot{\boldsymbol{\zeta}}+\mathbf{C^{T}}\ddot{\boldsymbol{\zeta}^{\prime}}+\mathbf{B}\ddot{\boldsymbol{\zeta}^{\prime}}+\dot{\mathbf{P}}\dot{\boldsymbol{\zeta}}+\dot{\mathbf{C^{T}}}\dot{\boldsymbol{\zeta}^{\prime}}+\dot{\mathbf{B}}\dot{\boldsymbol{\zeta}^{\prime}}=\mathbf{0},$ (20b)
@@
$\begin{split}\ddot{\boldsymbol{\zeta}}(l)&\equiv\mathbf{P}\ddot{\boldsymbol{\zeta}}+\mathbf{C^{T}}\ddot{\boldsymbol{\zeta}^{\prime}}+\mathbf{B}\ddot{\boldsymbol{\zeta}^{\prime}}+\dot{\mathbf{P}}\dot{\boldsymbol{\zeta}}+\dot{\mathbf{C^{T}}}\dot{\boldsymbol{\zeta}^{\prime}}+\dot{\mathbf{B}}\dot{\boldsymbol{\zeta}^{\prime}}\\ &+\frac{\partial^{3}B}{\partial\boldsymbol{\zeta}\partial\xi_{l}^{2}}\dot{\xi}_{l}^{2}+\frac{\partial^{3}B}{\partial\boldsymbol{\zeta}^{2}\partial\xi_{l}}\dot{\boldsymbol{\zeta}}\dot{\xi}_{l}^{2}+\frac{\partial^{2}B}{\partial\boldsymbol{\zeta}\partial\xi_{l}}\ddot{\xi}_{l}=\mathbf{0},\end{split}$ (21b)
Consequently, the term $-\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ can be rewritten using the identity given by (13)) as
@@
From this point, we separate the analyses for the cases of varying parameter at the fixed end and the free end.

### v5

@@ 1 Introduction
The study of physical problems using the variational framework has a rich history. In this setting, the equilibria are characterized as critical points of an energy-like functional. Typically, these problems are expressed in terms of a control parameter, generating a parameter family of equilibria. Indeed, the equilibria are computed by a slow scan of a parameter from a known solution using a continuation technique [1]. However, not all equilibria are stable and physically relevant; they must correspond to the local minima of the functional. Hence, it is crucial to determine whether the critical points are local minima (or stable) and track their changes. Generally, the stability of parameter-dependent equilibria changes at singularities, such as a fold or bifurcation in the parameter space [2, 3, 4]. However, it remains unclear whether a fold in the unstable equilibrium results in a stable equilibrium or a higher-order unstable equilibrium. The direction of stability transitions can be encoded through a particular projection of solutions, using distinguished bifurcation diagrams, and the stability can be assessed with a little observation and without any additional analysis. This idea of topologically exploiting the variational structure for stability examination was first employed in Astrophysics [5, 6]. Distinguished bifurcation diagrams were later developed for finite-dimensional mechanical systems [7] that are subsequently extended to infinite-dimensional systems [8, 9]. Since then, several versions of these diagrams have appeared in investigations of elastic problems [10, 11]. In most studies, the diagrams are limited to cases with fixed-fixed ends or cases where the varying parameter appears within the integrand or as a Lagrange multiplier. In this study, we generalize the distinguished bifurcation diagrams to variational problems subject to fixed-free ends, when the bifurcation parameter appears in the boundary conditions. Consequently, two relevant cases arise: when the parameter appears at the fixed end and when it appears at the free end. For this analysis, we introduce the notion of a stability index that is equal to the dimension of the subspace over which the associated second variation operator is negative. The critical points with index zero are local minima and stable. Alternatively, the energy of the system can be used to determine the direction of stability [12, 13]. The lower branch of the fold corresponds to the lower-energy state and therefore, local minima.
@@
The developed framework is applied to test the stability of elastica under an end load through a lever arm. Cantilever structures with a lever arm are common in nature and engineering—for example, in plants supporting fruits or in load-bearing flexible systems. In recent years, nonlinear rod models [14] have been actively employed in the fields of biophysics [15], computer graphics [16, 17] and soft robot manipulators [18, 19]. Elastica, a simple planar version of this rod model, has equally received attention. A flexible elastic structure subjected to loads or constraints may exhibit multiple equilibria. Hence, it is crucial to determine their stability, as only stable structures are physically viable. Unstable equilibria may exist under specific constraints or boundary conditions and transitions to the available stable equilibrium under disturbances. Indeed, the presence of instability also hints at snap-back phenomena, a characteristic of multi-stable systems [20]. Numerous studies have been conducted on the stability of elastic rods, which involve either the Jacobi test [21, 22, 23, 24, 25, 26] or Hessian eigenvalue determination [27, 28]. Alternative techniques based on bifurcation analysis are also available [29]. In this article, we assess the stability without performing any such rigorous computations. But, instead, conclusions are drawn by qualitatively examining the plot of appropriate projection of the equilibria against the parameter.
@@
Advances in material science produced highly deformable alloys and polymers, accelerating the development of soft robotics. Generally, soft robots employ very flexible structures to generate compliant mechanisms. A flexible elastic structure under load can exhibit multiple equilibria, where one equilibrium may transition to another through snap-back instability under disturbances. These bistable or multi-stable systems exhibit large rapid displacements for small stimuli when operated around the regions of instabilities. As the system surpasses the initial energy barrier, it transitions to a new state, accompanied by a release of energy. Some of nature’s quickest mechanisms seen in Mantis shrimp [30], Venus fly trap [31], and Hummingbird beak [32] can be attributed to this snap-back instability. These mechanisms are widely used in the design of high-performance soft actuators [33], soft robots [34], deployable and morphing structures [35] and haptic devices [36]. Inspired by these expanding applications, we aim to investigate elastica with a lever arm and the dependence of its stability on system parameters. These findings could offer invaluable insights for designing innovative mechanisms. Particularly, this study signifies the distinguished bifurcation diagrams for the analysis of multi-stable systems in engineering applications.
@@
This article is organized as follows. The section 2 introduces the unconstrained calculus of variation problem with a parameter in the boundary conditions and describes the associated distinguished bifurcation diagrams. Section 3.1 formulates the elastica with an end load acting through a lever arm. A bifurcation analysis of resulting equilibria is also performed. In section 4, we present several examples investigating the impact of varying parameters on stability using distinguished bifurcation diagrams. Finally, section 5 provides a summary and discussion of the results.
@@ 2 The Unconstrained Variational Problem
In this section, we extend the theory of distinguished bifurcation diagrams developed by [8] and [9] to variational problems subject to fixed-free ends. We consider both cases, where the bifurcation parameter appears at the fixed end and when it appears at the free end. In section 2.1, the problem and the relevant notation are introduced. The distinguished diagram at a simple fold when the parameter appears at the fixed end is developed in section 2.2 and when the parameter appears at the free end is developed in section 2.3.
@@ 2.1 Problem Formulation
Let $\boldsymbol{\zeta}:s\rightarrow\mathbb{R}^{p}$ be a vector-valued continuous differentiable function for $s\in[0,l]$. Consider an unconstrained calculus of variations problem of the form
$J(\boldsymbol{\zeta},\xi_{0},\xi_{l})=\int_{0}^{l}\mathcal{L}(\boldsymbol{\zeta},\boldsymbol{\zeta}^{\prime},s)ds+B(\boldsymbol{\zeta},\xi_{l},l),$ (1)
@@
while the other end $s=l$ is held free. The scalar function $\mathcal{L}$ has continuous second derivatives with respect to all its arguments and is convex in its second argument. On the other hand, the scalar function $B$ has continuous second derivatives with respect to its first argument and has continuous first derivatives with respect to the other arguments. The system can be influenced by the parameters $\xi_{0}$ and $\xi_{l}$ at the boundaries $s=0$ and $s=l$ respectively. In the current study, we vary either $\xi_{0}$ or $\xi_{l}$, but not both at a time. Maddocks [8] generated the bifurcation diagrams for problems when the parameter appears in the integrand and with homogeneous boundary conditions, which were later extended to the case when the parameter appears in the fixed end in a problem with non-homogeneous boundary conditions [9]. However, these studies are confined to cases of fixed-fixed ends. Here, we extend the analysis to the case where one of the fixed ends is set free. The first-order condition yields the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, $i=0$ or $l$, of (1) as solutions to the well-known Euler-Lagrange equations
@@
The notation $\xi_{i}$ refers to either $\xi_{0}$ or $\xi_{l}$ when not explicitly specified. These critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$ exist as a continuous curve of solutions along a branch for some pseudo-arclength parameterization $\tau$ [1] satisfying the property
$\big{|}\dot{\boldsymbol{\zeta}}(s,\tau)\big{|}^{2}+\big{|}\dot{\xi_{i}}(\tau)\big{|}^{2}=1.$
The notation $\dot{\boldsymbol{\zeta}}(s,\tau)$ denotes the derivative with respect to pseudo-arclength $\tau$, whereas ${\boldsymbol{\zeta}}^{\prime}(s,\tau)$ denotes the derivative with respect to independent parameter $s$. The second variation $\delta^{2}J$ evaluated at the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, must satisfy the condition [37]:
$\begin{split}\delta^{2}J(\boldsymbol{\zeta}_{o})[\mathbf{h}]=\frac{1}{2}&\int_{0}^{l}\left(\mathbf{h}^{\prime}\cdot\mathbf{P}\mathbf{h}^{\prime}+\mathbf{h}\cdot\mathbf{C}\mathbf{h}^{\prime}+\mathbf{h}^{\prime}\cdot\mathbf{C}^{T}\mathbf{h}+\mathbf{h}\cdot\mathbf{Q}\mathbf{h}\right)ds+\frac{1}{2}\mathbf{B}\mathbf{h}(l)\cdot\mathbf{h}(l)\geq 0,\end{split}$ (5)
@@
For brevity, the dependence of $\mathbf{P},\mathbf{C}$ and $\mathbf{Q}$ on $s$ is not explicitly shown. The matrices $\mathbf{P}$$\mathbf{Q}$ and $\mathbf{B}$ are symmetric, whereas the matrix $\mathbf{C}$ may not be. Here, $\mathbf{h}$ is a variation in the solution that satisfy the linearized boundary condition:
@@
where $\mathcal{S}$ is the second-order differential operator
@@
and $\langle\cdot,\cdot\rangle$ represents the standard inner-product in $L^{2}$- space. The coefficient $\frac{1}{2}$ has no effect on the results and is omitted in the subsequent analysis. We define an notion of index [38] which denotes the dimension of the $\mathbf{h}$ subspace over which the second variation denoted by (9) is negative. The zero index solutions satisfy the necessary condition for local minima given by (5) and correspond to stable solutions. In fact, the index can be interpreted as the number of eigenvalues of the eigenvalue problem
@@
The operator $\mathcal{S}$ is assumed to have eigenvalues and eigenfunctions that smoothly depend on the pseudo-arclength $\tau$. Therefore, it is not sufficient to merely determine the sign of the eigenvalue; it is also necessary to account for the additional boundary contribution. Alternatively, one can consider the (5) directly using numerical methods. Upon discretization, it leads to a Global Stiffness Matrix whose eigenvalues dictate stability - positive eigenvalues imply local minimum and thus indicate stability. The index for the case of fixed-fixed ends is shown to be equal to the number of conjugate points (for more details refer to [21]). However, we will not pursue our analysis in that direction. Instead, we track changes in the index at critical points.
@@
$\langle\mathcal{S}\mathbf{u},\mathbf{v}\rangle=\langle\mathbf{u},\mathcal{S}\mathbf{v}\rangle+\Big{[}\left(\mathbf{P}\mathbf{v}^{\prime}+\mathbf{C}^{T}\mathbf{v}\right)\cdot\mathbf{u}\Big{]}^{l}_{0}-\Big{[}\left(\mathbf{P}\mathbf{u}^{\prime}+\mathbf{C}^{T}\mathbf{u}\right)\cdot\mathbf{v}\Big{]}^{l}_{0},$ (13)
through successive applications of integration by parts. Therefore, when both the boundary terms vanish $\mathcal{S}$ is a self-adjoint operator.
Parameter-dependent variational problems generate a family of extremals, and in the case of a single parameter $\xi$, a curve of solutions is obtained. Sometimes, these solutions are characterized by folds- points at which the curve slopes vertically. In bifurcation terminology, these points are also referred to as saddle-node bifurcation or turning points. On differentiating the Euler-Lagrange equations given by (3) and the non-linear boundary conditions given by (4) with respect to pseudo-arclength $\tau$, we obtain the following expressions for the two cases of varying parameters
@@
The Hessian matrices that appear in the boundary terms $s=l$ are evaluated at that boundary, and their dependence on $l$ is not included for conciseness. Each of these boundary value problems corresponds to the eigenvalue problem denoted by (11) with zero eigenvalue whenever $\dot{\xi}_{0}=0$ or $\dot{\xi}_{l}=0$, with the eigenvector $\mathbf{h}$ coinciding with $\dot{\boldsymbol{\zeta}}$. The boundary condition at $s=l$ also removes the boundary term in (12), fetching a zero second-variation quadratic functional. We focus on the behavior of the motion of this functional around these special points - folds. A simple fold further satisfies:
$\ddot{\xi}_{i}\neq 0,\qquad$ (15a)
$\frac{d}{d\tau}J_{\xi_{i}}\neq 0,\qquad$ (15b)
$\text{and }\mu=0\text{ is a simple eigenvalue of }\mathcal{S}.$ (15c)
An eigenvalue is called a simple eigenvalue if its algebraic multiplicity is one. The meaning of (15b) becomes clearer towards the end. We analyze the behavior of the second-variational functional in the vicinity of these folds, which involves simultaneously monitoring the eigenvalues and the boundary term contribution from the corresponding eigenfunctions. We assume that changes in stability occur exclusively in the vicinity of critical points, such as folds or bifurcation points [2].
@@
$\frac{d}{d\tau}\delta^{2}J[\mathbf{h}]=\langle\dot{\left(\mathcal{S}\mathbf{h}\right)},\mathbf{h}\rangle+\dot{\bigg{(}\mathbf{P}\mathbf{h}^{\prime}+\mathbf{C}^{T}\mathbf{h}+\mathbf{B}\mathbf{h}\bigg{)}}\cdot\mathbf{h}\bigg{|}_{s=l}+\langle\mathcal{S}\mathbf{h},\dot{\mathbf{h}}\rangle+\bigg{(}\mathbf{P}\mathbf{h}^{\prime}+\mathbf{C}^{T}\mathbf{h}+\mathbf{B}\mathbf{h}\bigg{)}\cdot\dot{\mathbf{h}}\bigg{|}_{s=l},$ (16)
@@
Since $\dot{\boldsymbol{\zeta}}$ is the zero eigenvector of $\mathcal{S}$, we have $\mathcal{S}\dot{\boldsymbol{\zeta}}=\mathbf{0}$ implying $\langle\dot{\mathbf{h}},\mathcal{S}\dot{\boldsymbol{\zeta}}\rangle=0$. On applying the boundary conditions denoted by (14) and (17) at a fold $\dot{\xi_{i}}$, the above expression simplifies to
@@
$\mathbf{P}\ddot{\boldsymbol{\zeta}}(l)$ $+\mathbf{C}^{T}\ddot{\boldsymbol{\zeta}^{\prime}}(l)+\mathbf{B}\ddot{\boldsymbol{\zeta}^{\prime}}(l)+\dot{\mathbf{P}}\dot{\boldsymbol{\zeta}}(l)+\dot{\mathbf{C}^{T}}\dot{\boldsymbol{\zeta}^{\prime}}(l)+\dot{\mathbf{B}}\dot{\boldsymbol{\zeta}^{\prime}}(l)=\mathbf{0},$ (20b)
@@
$\begin{split}\mathbf{P}\ddot{\boldsymbol{\zeta}}(l)&+\mathbf{C}^{T}\ddot{\boldsymbol{\zeta}^{\prime}}(l)+\mathbf{B}\ddot{\boldsymbol{\zeta}^{\prime}}(l)+\dot{\mathbf{P}}\dot{\boldsymbol{\zeta}}(l)+\dot{\mathbf{C}}^{T}\dot{\boldsymbol{\zeta}^{\prime}}(l)+\dot{\mathbf{B}}\dot{\boldsymbol{\zeta}^{\prime}}(l)\\ &+\frac{\partial^{3}B}{\partial\boldsymbol{\zeta}\partial\xi_{l}^{2}}\dot{\xi}_{l}^{2}+\frac{\partial^{3}B}{\partial\boldsymbol{\zeta}^{2}\partial\xi_{l}}\dot{\boldsymbol{\zeta}}(l)\dot{\xi}_{l}^{2}+\frac{\partial^{2}B}{\partial\boldsymbol{\zeta}\partial\xi_{l}}\ddot{\xi}_{l}=\mathbf{0},\end{split}$ (21b)
Consequently, the term $\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ can be rewritten using the identity given by (13)) as
@@
From this point, we separate the analyses for the cases of varying parameter at the fixed end $(\xi_{0})$ and the free end $(\xi_{l})$.